OpenStudy (anonymous):

solve by the substitution method 5x-2y=-37 6x+57=y

OpenStudy (anonymous):

pls post in math in math session 5x-2(6x+57)=-37 5x-12x-114=-37 -7x=77 x=-11 y=-66+57 y=-9